# Field equation of thermodynamic gravity and galactic rotational curves

### v2

@@ Abstract.
The rotational velocity curve (RC) of galaxy NGC 3198 is modelled in various theoretical frameworks: Thermodynamic Gravity (TG) is compared to Dark Matter (DM) and Modified Newtonian Dynamics (MOND). The nonlinear gravitational field equation of TG is solved using the baryonic mass density as the source of the gravitational field. In this paper, first a dissipation motivated numerical method is verified with the help of exact solutions. Then, the obtained optimal velocity curve is compared to pseudo-isothermal DM and MOND parametrisations for the RC of the galaxy.
@@ 1. Introduction
The most successful model in cosmology so far is the so-called $\Lambda$CDM model, where $\Lambda$ stands for the cosmological constant in the Einstein-equation accounting for dark energy, and CDM stands for cold dark matter. The composition of DM is unknown. There is an ongoing theoretical, experimental and computational hunt to find proper candidates in the framework of existing or new particles of the Standard Model of Particle Physics [1, 2, 3]. The most well-known effect of dark matter is its influence on the rotation curves of galaxies. Properly modelling the observed galactic rotational velocity curves with DM requires particular dark matter distributions, [4, 1]. However, in spite of the flexibility in modelling with the help of a potentially unknown distribution, there are some well-known problems and inconsistencies both in the case of individual galaxies and general, global properties. Regularities in baryon-DM coupling can be summarised in the Radial Acceleration Relation (RAR). This relation generalises the more well-known dynamical relations, like Renzo’s rule or the baryonic Tully-Fisher relation. The RAR encompasses these effects for various types of galaxies and establishes the strong connection between the baryonic contribution and the rotation curve [5, 6]. These regularities indicate that DM is not a simple passive background but interacts with the baryonic mass component. Related well-documented problems of the $\Lambda$CDM paradigm are the core-cusp problem, the missing satellite problem and the too-big-to-fail problem. The simulations that include dark matter predict that DM halo density profiles should rise steeply at small radii as $\rho(r)\propto r^{-\gamma}$ with $\gamma\simeq 0.8-1.4$ for small galaxies. This is in contrast to observations with many low-mass DM-dominated galaxies. A related ’central density problem’ also exists as simulations predict more dark matter in the central regions than they should host. These two issues (low-density cores and high-density cusps) are summarised as the core-cusp problem [7]. That is one of the reasons that the hydrodynamics-motivated pseudo-isothermal (ISO) DM distribution is widely used for rotational velocity curve modelling. Because of the above-mentioned problematic aspects, there are several other proposals to explain the missing mass. The superfluid theories suggest a particular multi-component form of dark matter, [8, 9, 10], where the explanation of rotational curves may come from a logarithmic field potential, [11, 12], or scale invariance, [13]. Among these suggestions, the most popular and most developed is MOND. Its success is based on the insight regarding the nonrelativistic, phenomenological observations, in particular, the galactic rotational curves. The reason became clear after the development of the first modified gravity version of MOND, called A QUAdratic Lagrangian theory (AQUAL), [14]. It has a vacuum solution with a force that is inversely proportional to the distance of the central mass: that can exactly counterbalance the centrifugal force and explain the typical flat velocity profiles.
@@
A different approach was proposed by Verlinde, whose emergent gravity (also known as entropic gravity) uses the idea that spacetime and gravity emerge together from the entanglement structure of an underlying microscopic theory. Using insights from black hole thermodynamics, quantum information theory and string theory, emergent gravity is motivated and borrows elements of various nonrelativistic constitutive theories. For the DM phenomena, emergent gravity is motivated by generalised nonrelativistic elasticity [17]. The AQUAL type field equation in the nonrelativistic limit can be obtained from a covariant version of the theory, [18]. There is also a theory of gravity based on nonequilibrium thermodynamics, refered as thermodynamic gravity (TG) in the following, [19, 20]. Then the gravitational potential is treated as a thermodynamic state variable and the standard Newtonian gravity can be derived. The approach is based on the simple and only assumption that field energy and interaction energy is included in the total energy density. Then the inequality describing the entropy production rate in a local equilibrium state leads to the following dissipative field equation for gravity:
@@
where $\varphi$ is the gravitational potential field, $l$ refers to the macroscopic spatial scale of variations of the whole system, $\tau$ is the relaxation time, $\Delta$ is the Laplace operator, $K$ is a characteristic coefficient of the dimensions $s^{2}m^{-2}$, $G$ is the Newtonian constant of gravity and $\rho$ is the mass density. The stationary solution yields a nonlinear field equation for a modified gravitational interaction, while the strength of deviation from the classical case is characterised by the $K$ constant, which arises from the linear coupling between the gravitational and mechanical interactions. This leads to a dynamic crossover between Newtonian and non-Newtonian regimes. There, the relaxation to a stationary nonlinear field equation indicates the dissipative character of the theory. The nonlinear term is due to the presence of the gravitational field pressure, which is a consequence of the holographic properties of Newtonian gravity, and in general, the result of thermodynamic requirements, [21]. Therefore, due to the pressure of the gravitational field, there is a cross effect between the thermodynamic force of the scalar gravitational interaction and the scalar part of the mechanical interaction and the nonlinear term of eq. 1 emerges naturally. Therefore, the parameter $K$ is not fundamental, it belongs to a particular transport property of the thermodynamic cross effect of gravitational field and mechanical momentum transport: the gravitational field transports material momentum and material momentum pulls the related gravity field along. $K$ is a material parameter that characterises the intensity of the aforementioned coupling in a galactic environment, therefore it is not a self-energy term as in the similar field equations of [22, 23, 24]. Also, it has a direct physical interpretation, as $v_{K}=K^{-1/2}$ is a characteristic velocity of a flat rotational curve because $F_{K}=\frac{1}{Kr}$ is the force field in a stationary vacuum solution of the field equation in TG. In the TG theory, the source term of the field equation is the density distribution of the baryonic mass, the sum of the visible matter and the atomic hydrogen densities. The gravitational field and the corresponding centrifugal force can be calculated by solving the stationary part of eq. 1. This paper serves as a proof of concept for this method, to be detailed for additional galaxies in following publications. In the following sections, we give a finite difference scheme to solve the field equations and verify the method with the help of known exact solutions of the equation with constant, $\rho_{0}\geq 0$ density distributions. Then, we demonstrate the applicability of the method, solving the equation with a realistic example, with the mass distribution of galaxy NGC 3198 as a source term. Here the sum of the visible part (referenced also as a stellar disk) and the atomic H (HI) component is used to solve the stationary part of the field equation 1. The numerical method is of a relaxational type, exploiting that the time-dependent field equation relaxes to the stationary solution. The paper first presents the exact solutions of the equation. Then, a short description of the numerical method is given. Afterwards, we validate the method, showing that the numerical scheme relaxes to the known exact solutions. Then, the velocity curve of the NGC 3198 galaxy is modelled, fitting the K parameter. Finally, a short analysis and interpretation follow.
@@ 2.1. Exact solutions
This non-linear Ricatti equation has well-known solutions when the density is zero or a constant value.
@@ 2.1.1. Vacuum solution
(14) $\varphi(r)=-\frac{1}{K}\ln{\frac{r}{K+Cr}}+\varphi_{0}.$
@@ 2.2.2. Discretisation
during the galaxy velocity curve calculation. For the reliability testing, the boundaries were set to the corresponding values of the theoretical curves.
@@ 3.1. Methodology
The velocities corresponding to the surface densities for atomic hydrogen (HI) and the stellar disk (SD) were converted into an equivalent pseudo-spherical dataset to calculate density data. The original velocities are defined as the velocities each component would induce in the plane of the galaxy, and for the comparison a spherical case is considered. The work by de Blok et al. [29] uses the convention that the negative sign of the HI data means that it has a negative contribution when calculating the total $v^{2}$ sum, that is, it has an outward gravitational effect due to the lack of spherical symmetry and the specifics of its distribution. The observed velocity curve data is based on the $21$-cm HI observations, the SD data is calculated from luminosity and the derived HI gas rotation curve is from the hydrogen column density. To calculate the $\rho_{j}$ for the numerical method from the interpolated function of the rotation velocities $v_{rot}(r)$, the following procedure was used:
(47) $\rho_{j}=\left[\frac{\partial(v_{rot,SD}(r)^{2}r)}{\partial r}\right]_{j}\cdot\frac{1}{4\pi Gr_{j}^{2}}+\left[\frac{\partial(v_{rot,HI}(r)^{2}r)}{\partial r}\right]_{j}\cdot\frac{1}{4\pi Gr_{j}^{2}},$
and $\rho_{d}$ was selected to be the mean of the array $\rho_{j}$:
@@
For comparison, the pseudo-isothermal (ISO) DM halo model is used, as in the work of de Blok et al. [29] and Randriamampandry and Carignan [30]. This model has the density profile
(50) $\rho_{\mathrm{ISO}}(r)=\rho_{0}\left[1+\left(\frac{r}{R_{C}}\right)^{2}\right]^{-1},$
where $\rho_{0}$ is the central density and $R_{C}$ is the core radius of the halo. The corresponding dark matter rotation curve is given by
(51) $v(r)=\sqrt{4\pi G\rho_{0}R_{C}^{2}\left[1-\frac{R_{C}}{r}\arctan\left(\frac{r}{R_{C}}\right)\right]}.$
@@
For the galactic calculation, the $\rho_{0}$ and $R_{C}$ parameters are taken from Table 3 of [29], which contains mass models with fixed mass-luminosity ratios ($Y^{3.6}_{*}$, denoting the $3.6$ $\mu$m band) and Diet-Salpeter stellar initial mass function. To obtain the velocity curve from the ISO DM model, the contributions of the SD and HI are added to this. Regarding the MOND model, the method and data of Randriamampandry and Carignan are used, [30], which can be interpreted as the vacuum AQUAL solution substituted directly into the velocity calculation. The exact form of the interpolating function of MOND is not fixed by the theory itself and some forms work better in explaining certain phenomena. Zhao & Famaey [31] found that a simplified form of the interpolating function not only provides good fits to the observed rotational curves but also the derived mass/luminosity ratios are more compatible with those obtained from stellar populations synthesis models. In [30] this simple interpolating function of MOND is applied:
(52) $\mu(x)=\frac{x}{1+x},$
and the fitted velocity curve is given by
(53) $v^{2}(r)=\sqrt{v_{SD}^{2}+v_{HI}^{2}}\cdot\sqrt{a_{0}\cdot r+v_{SD}^{2}+v_{HI}^{2}},$
with $v_{SD}$ describing the contribution of the stellar disk (and bulge, if given separately) and $v_{HI}$ describing the contribution of the neutral hydrogen. The acceleration parameter, $a_{0}$, is fitted to the observational data using eq. 53.
@@ 3.2. Galactic rotational velocity curve of NGC 3198
In this section, the above-mentioned method is applied to reproduce the velocity curve of NGC 3198.
@@
ISO Halo model
$\rho_{0}$ [$10^{-3}M_{\odot}pc^{-3}$] $14.4\pm 2.0$
$R_{C}$ [$kpc$] $4.97\pm 0.41$
MOND
$a_{0}$ [$10^{-10}ms^{-2}$] $0.67\pm 0.02$
@@
$\tilde{K}$ $45$
$K$ [$s^{2}km^{-2}$] $3.6\cdot 10^{-5}$
Figure 6. NGC 3198 galaxy, source: Sloan Digital Sky Survey.
Figure 7. TG, MOND and ISO DM parameters for galaxy NGC 3198 [29].
Figure 8. Velocity curves from thermodynamic gravity, MOND and ISO dark matter for the NGC 3198 galaxy.
Figure 9. Velocity curves from thermodynamic gravity for galaxy NGC 3198 with the stellar disk mass distribution alone as the source for the gravitational field and the previous case with atomic H gas considered as well.

### v3

@@ Abstract.
The rotational velocity curve (RC) of galaxy NGC 3198 is modelled in various theoretical frameworks: Thermodynamic Gravity (TG) is compared to Dark Matter (DM) and Modified Newtonian Dynamics (MOND). The nonlinear gravitational field equation of TG is solved using the baryonic mass density as the source of the gravitational field. In this paper, first, a dissipation motivated numerical method is verified with the help of exact solutions. Then, the obtained optimal velocity curve is compared to DC14 model DM and MOND EFE (Modified Newtonian Dynamics - External Field Effect aspect) parametrisations for the RC of the galaxy.
@@ 1. Introduction
The most successful model in cosmology so far is the so-called $\Lambda$CDM model, where $\Lambda$ stands for the cosmological constant in the Einstein-equation accounting for dark energy, and CDM stands for cold dark matter. The composition of DM is unknown. There is an ongoing theoretical, experimental and computational hunt to find proper candidates in the framework of existing or new particles of the Standard Model of Particle Physics [1, 2, 3].
The most well-known effect of dark matter is its influence on the rotation curves of galaxies. Properly modelling the observed galactic rotational velocity curves with DM requires particular dark matter distributions, [4, 1, 5, 6]. However, in spite of the flexibility in modelling with the help of a potentially unknown distribution, there are some well-known problems and inconsistencies both in the case of individual galaxies and general, global properties.
Regularities in baryon-DM coupling can be summarised in the Radial Acceleration Relation (RAR). This relation generalises the more well-known dynamical relations, like Renzo’s rule or the baryonic Tully-Fisher relation. The RAR encompasses these effects for various types of galaxies and establishes the strong connection between the baryonic contribution and the rotation curve [7, 8]. These regularities indicate that DM is not a simple passive background but interacts with the baryonic mass component.
Related well-documented problems of the $\Lambda$CDM paradigm are the core-cusp problem, the missing satellite problem and the too-big-to-fail problem. The simulations that include dark matter predict that DM halo density profiles should rise steeply at small radii as $\rho(r)\propto r^{-\gamma}$ with $\gamma\simeq 0.8-1.4$ for small galaxies. This is in contrast to observations with many low-mass DM-dominated galaxies. A related ’central density problem’ also exists as simulations predict more dark matter in the central regions than they should host. These two issues (low-density cores and high-density cusps) are summarised as the core-cusp problem [9, 10]. That is one of the reasons that hydrodynamics-motivated DM distributions are widely used for rotational velocity curve modelling.
Because of the above-mentioned problematic aspects, there are several other proposals to explain the missing mass. The superfluid theories suggest a particular multi-component form of dark matter, [11, 12, 13, 6, 14], where the explanation of rotational curves may come from a logarithmic field potential, [15, 16], or scale invariance, [17].
Among these suggestions, the most popular and most developed is MOND. Its success is based on the insight regarding the nonrelativistic, phenomenological observations, in particular, the galactic rotational curves. The reason became clear after the development of the first modified gravity version of MOND, called A QUAdratic Lagrangian theory (AQUAL), [18]. It has a vacuum solution with a force that is inversely proportional to the distance of the central mass: that can exactly counterbalance the centrifugal force and explain the typical flat velocity profiles.
@@
A different approach was proposed by Verlinde, whose emergent gravity (also known as entropic gravity) uses the idea that spacetime and gravity emerge together from the entanglement structure of an underlying microscopic theory. Using insights from black hole thermodynamics, quantum information theory and string theory, emergent gravity is motivated and borrows elements of various nonrelativistic constitutive theories. For the DM phenomena, emergent gravity is motivated by generalised nonrelativistic elasticity [21]. The AQUAL type field equation in the nonrelativistic limit can be obtained from a covariant version of the theory, [22].
There is also a theory of gravity based on nonequilibrium thermodynamics, refered as thermodynamic gravity (TG) in the following, [23, 24]. Then the gravitational potential is treated as a thermodynamic state variable and the standard Newtonian gravity can be derived. The approach is based on the simple and only assumption that field energy and interaction energy is included in the total energy density. Then the inequality describing the entropy production rate in a local equilibrium state leads to the following dissipative field equation for gravity:
@@
where $\varphi$ is the gravitational potential field, $l$ refers to the macroscopic spatial scale of variations of the whole system, $\tau$ is the relaxation time, $\Delta$ is the Laplace operator, $K$ is a characteristic coefficient of the dimensions $s^{2}m^{-2}$, $G$ is the Newtonian constant of gravity and $\rho$ is the mass density. The stationary solution yields a nonlinear field equation for a modified gravitational interaction, while the strength of deviation from the classical case is characterised by the $K$ constant, which arises from the linear coupling between the gravitational and mechanical interactions. This leads to a dynamic crossover between Newtonian and non-Newtonian regimes in the vacuum region.
There, the relaxation to a stationary nonlinear field equation indicates the dissipative character of the theory. The nonlinear term is due to the presence of the gravitational field pressure, which is a consequence of the holographic properties of Newtonian gravity, and in general, the result of thermodynamic requirements, [25, 26].
Therefore, due to the pressure of the gravitational field, there is a cross effect between the thermodynamic force of the scalar gravitational interaction and the scalar part of the mechanical interaction and the nonlinear term of eq. 1 emerges naturally. Therefore, the parameter $K$ is not fundamental, it belongs to a particular transport property of the thermodynamic cross effect of gravitational field and mechanical momentum transport: the gravitational field transports material momentum and material momentum pulls the related gravity field along. $K$ is a material parameter that characterises the intensity of the aforementioned coupling in a galactic environment, therefore it is not a self-energy term as in the similar field equations of [27, 28, 29]. Also, it has a direct physical interpretation, as $v_{K}=K^{-1/2}$ is a characteristic velocity of a flat rotational curve because $F_{K}=\frac{1}{Kr}$ is the force field in a stationary vacuum solution of the field equation in TG.
In the TG theory, the source term of the field equation is the density distribution of the baryonic mass, the sum of the visible matter and the atomic hydrogen densities. The gravitational field and the corresponding centrifugal force can be calculated by solving the stationary part of eq. 1. This paper serves as a proof of concept for this method, to be detailed for additional galaxies in following publications.
In the following sections, we give a finite difference scheme to solve the field equations and verify the method with the help of known exact solutions of the equation with constant, $\rho_{0}\geq 0$ density distributions. Then, we demonstrate the applicability of the method, solving the equation with a realistic example, with the mass distribution of galaxy NGC 3198 as a source term. Here the sum of the visible part (referenced also as a stellar disk) and the atomic H (HI) component is used to solve the stationary part of the field equation 1. The numerical method is of a relaxational type, exploiting that the time-dependent field equation relaxes to the stationary solution.
The paper first presents the exact solutions of the equation. Then, a short description of the numerical method is given. Afterwards, we validate the approach, showing that the numerical scheme relaxes to the known exact solutions. Then, the velocity curve of the NGC 3198 galaxy is modelled, fitting the K parameter. Finally, a short analysis and interpretation follow.
@@ 2.1. Exact solutions
This non-linear Riccati equation has well-known solutions when the density is zero or a constant value.
@@ 2.1.1. Vacuum solution
(14) $\varphi(r)=\frac{1}{K}\ln{\frac{r}{K+Cr}}+\varphi_{0}.$
@@ 2.2.2. Discretisation
during the galaxy velocity curve calculation, with $r_{min}$ denoting the radial position of the innermost observed velocity data point. For the reliability testing, the boundaries were set to the corresponding values of the theoretical curves.
@@ 3.1. Methodology
[…]
with $\Upsilon_{1}=M_{*}/L$ denoting the fitted mass-luminosity ratio and $\rho_{d}$ was selected to be the mean of the array $\rho_{j}$:
@@
For comparison, we show a dark matter and a MOND modelling parametrisation. The dark matter model with DC14 density profile and with flat prior is used, from the work of Li et al. [35, 36]. This model has the density profile
(50) $\rho_{\mathrm{DC14}}(r)=\frac{\rho_{s}}{\left(\frac{r}{r_{s}}\right)^{\gamma}\left[1+\left(\frac{r}{r_{s}}\right)^{\alpha}\right]^{(\beta-\gamma)/\alpha}},$
where $\rho_{s}$ is a characteristic volume density, $r_{s}$ is the scale radius; $\beta$ and $\gamma$ are the inner and outer slopes respectively, and $\alpha$ describes the transition between the inner and outer regions. The last three paramaters’ value depends on the stellar-to halo mass ratio (adopted from [35]), and ultimately, the corresponding dark matter rotation curve is given by
(51) $v_{\mathrm{DC14}}(r)=v_{200}\sqrt{\frac{C_{200}}{x}\frac{B(a,b+1,\epsilon)+B(a+1,b,\epsilon)}{B(a,b+1,\epsilon_{c})+B(a+1,b,\epsilon_{c})}},$
with $a=(3-\gamma)/\alpha$, $b=(\beta-3)/\alpha$, $\epsilon=\frac{(r/r_{s})^{\alpha}}{1+(r/r_{s})^{\alpha}}$, $\epsilon_{c}=\frac{C_{200}^{\alpha}}{1+C_{200}^{\alpha}}$, $B(a,b,x)$ being the incomplete beta-function and $x=\frac{r}{r_{s}}$. The above formula is obtained after calculating the total mass of the DC14 mass distribution inside a given radius $r$. The $C_{200}$ and $v_{200}$ parameters were used in their fitting method, and the relate to the fitted physical parameters as
(52) $C_{200}=r_{200}/r_{s},\leavevmode\nobreak\ v_{200}=10C_{200}r_{s}H_{0},$
with $r_{200}$ being the virial radius inside of which the average density is 200 times the critical density of the Universe ($\rho_{\mathrm{crit}}=\frac{3H_{0}^{2}}{8\pi G}$) and $H_{0}=73\mathrm{kms}^{-1}\mathrm{Mpc}^{-1}$ is the Hubble constant. For the galactic calculation, the fitted parameters for the flat prior DC14 dark matter model are taken from the SPARC website [^1] http://astroweb.cwru.edu/SPARC from [35, 36]. To obtain the velocity curve from the DC14 DM model, the contributions of the SD and HI are added to this:
(53) $v_{\mathrm{DM}}(r)=\sqrt{v_{\mathrm{DC14}}^{2}(r)+(M_{*}/L)v_{rot,SD}(r)^{2}+v_{rot,HI}(r)\cdot\mathrm{abs}\left({v_{rot,HI}(r)}\right)}.$
@@
Regarding MOND, the method and data of Chae et al. are used, [37]. The exact form of the interpolating function of MOND is not fixed by the theory itself and some forms work better in explaining certain phenomena. Zhao & Famaey [38] found that a simplified form of the interpolating function not only provides good fits to the observed rotational curves but also the derived mass/luminosity ratios are more compatible with those obtained from stellar populations synthesis models. It is also important to remark, that considering the observation uncertainties, like with the mass/luminosity ratio, not only the quality of the modelling is improved, but also there is no need to deviate from the standard MOND acceleration, $a_{0}$, contrary to the modelling approach of Wang and Chen [39]. In [37] the expected acceleration of a test particle ($v^{2}/r$ for circular motion) is given by
(54) $g=\nu_{e_{N}}\left(y\equiv g_{N}/a_{0}\right)g_{N},$
with $g_{N}$ as the Newtonian internal gravitational field (from the galactic matter), $a_{0}=1.2\cdot 10^{-10}\mathrm{ms}^{-2}$ MOND acceleration scale and $\nu_{e_{N}}$ being
(55) $\nu_{e_{N}}(y)=\frac{1}{2}+\sqrt{\frac{1}{4}D^{2}_{e_{N}}(y)+\frac{D_{e_{N}}(y)}{y}}-\frac{e_{N}}{\sqrt{|e_{N}|}}\frac{C_{e_{N}}}{y},$
where $C_{e_{N}}\equiv\sqrt{1+|e_{N}|/4}$, $D_{e_{N}}(y)\equiv 1+|e_{N}|/y$ and the fitted parameter is $\tilde{e}\equiv e_{N}/\sqrt{|e_{N}|}$. Here, $g_{N}$ is given as
(56) $g_{N}=\frac{v_{\mathrm{bar}}^{2}}{r},$
with
(57) $v_{\mathrm{bar}}(r)=\sqrt{(M_{*,\mathrm{disk}}/L)v_{rot,SD}(r)^{2}+(M_{\mathrm{gas}}/M_{\mathrm{HI}})v_{rot,HI}(r)\cdot\mathrm{abs}\left({v_{rot,HI}(r)}\right)}$
describing the contribution of the stellar disk and the neutral hydrogen. Both the DM and MOND EFE methods fit the distance to the galaxy and its inclination, which can be summarised as
(58) $v_{\mathrm{bar,fit}}^{2}=\frac{D_{\mathrm{fit}}}{D_{\mathrm{obs}}}v_{\mathrm{bar}}^{2},$
(59) $v_{\mathrm{obs,fit}}=v_{\mathrm{obs}}\frac{\sin{i_{\mathrm{obs}}}}{\sin{i_{\mathrm{fit}}}},$
(60) $r_{\mathrm{fit}}=\frac{D_{\mathrm{fit}}}{D_{\mathrm{obs}}}r_{\mathrm{obs}}$
with $v_{\mathrm{obs}}$ denoting the observed rotational velocity of the galaxy and $v_{\mathrm{obs,fit}}$ its fit-adjusted value. In our method, this gives diminishing return, thus for the better quality of the algorithmic fit, we omit these parameters, but note that their uncertainties are important to consider when evaluating galactic RC fits.
@@ 3.2. Galactic rotational velocity curve of NGC 3198
In this section, the above-mentioned method is applied to reproduce the velocity curve of NGC 3198. The parameters $\Upsilon_{1}$ and $K$ were fitted using scipy and their presented error is the resulting estimate from the fit. The reduced chi-squared value is calculated as the following:
(61) $\chi^{2}_{\nu}=\sum_{r}\frac{\left(v_{obs}-v_{fit}\right)^{2}/\sigma_{obs}^{2}}{N-f},$
with $\sigma_{obs}$ denoting the uncertainties in the observed velocity curve, $N$ the number of data points and $N-f$ the degrees of freedom. In our case, $f=4$, as the two ends of the data set are used as fixed boundary points and two parameters are fitted. Due to our simpler, spherical model, our mass-luminosity ratio is not strictly physical (as a cylindrical model is usually preferred), but gives a plausible conversion factor for the fit.
[…]
@@
DC14 DM model
$\log(\rho_{s})$ [$\rho_{s}$: $M_{\odot}pc^{-3}$] $-1.46\pm 0.21$
$D$ [Mpc] $10.47\pm 1.17$
$r_{s}$ [$kpc$] $5.77\pm 1.23$
$i$ [deg] $71.5\pm 3.0$
$M_{*}/L$ $0.40\pm 0.09$
$\chi^{2}_{\nu}$ $1.26$
MOND EFE
$\tilde{e}$ $0.055^{+0.012}_{-0.011}$
$D$ [Mpc] $15.27^{+1.17}_{-1.08}$
$i$ [deg] $75.62^{+2.73}_{-2.71}$
$M_{\mathrm{gas}}/M_{\mathrm{HI}}$ $1.36^{+0.13}_{-0.12}$
$M_{*,\mathrm{disk}}/L$ $0.43^{+0.04}_{-0.04}$
$\chi^{2}_{\nu}$ $1.62$
@@
$K$ [$10^{-5}s^{2}km^{-2}$] $3.40\pm 0.04$
$M_{*}/L$ $0.762\pm 0.007$
$\chi^{2}_{\nu}$ $1.29$
Figure 7. TG, MOND EFE and DC14 DM parameters for galaxy NGC 3198 [37, 35, 36] .
Figure 8. Velocity curves from thermodynamic gravity, MOND EFE from Chae et al. [37] and DC14 (flat prior) from [35, 36] dark matter for the NGC 3198 galaxy.